Net Present Value—Unequal Lives

Bunker Hill Mining Company has two competing proposals: a processing mill and an electric shovel. Both pieces of equipment have an initial investment of $681,948. The net cash flows estimated for the two proposals are as follows:

Net Cash Flow
Year      Processing Mill      Electric Shovel
1 $218,000 $273,000
2 194,000 253,000
3 194,000 233,000
4 155,000 240,000
5 118,000
6 98,000
7 85,000
8 85,000

The estimated residual value of the processing mill at the end of Year 4 is $270,000.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Determine which equipment should be favored, comparing the net present values of the two proposals and assuming a minimum rate of return of 12%. Use the present value table appearing above.

Processing Mill Electric Shovel
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $
Answer #1
Calculation Of Present Value Of Cash Flow
Year Particular Discounting Factor @12% Processing Mill Electric Shovel
Cash Flow DCF Cash Flow DCF
a b c d e=d*c f g=f*c
1 Annual Cash Inflows 0.893 $            2,18,000 $               1,94,674 $          2,73,000 $         2,43,789
2 Annual Cash Inflows 0.797 $            1,94,000 $               1,54,618 $          2,53,000 $         2,01,641
3 Annual Cash Inflows 0.712 $            1,94,000 $               1,38,128 $          2,33,000 $         1,65,896
4 Annual Cash Inflows 0.636 $            1,55,000 $                   98,580 $          2,40,000 $         1,52,640
5 Annual Cash Inflows 0.567 $            1,18,000 $                   66,906
6 Annual Cash Inflows 0.507 $                98,000 $                   49,686
7 Annual Cash Inflows 0.452 $                85,000 $                   38,420
8 Annual Cash Inflows 0.404 $                85,000 $                   34,340
Present Value Of Cash Flow $               7,75,352 $         7,63,966
Processing Mill Electric Shovel
Present value of net cash flow total $                           7,75,352 $            7,63,966
Less amount to be invested 681948 681948
Net present value $                               93,404 $                82,018
Processing Mill equipment should be favored.